# Pan-cancer Histopathology WSI Pre-training with Position-aware Masked Autoencoder

### v2

@@ 4.2 Experimental setting
We first pre-trained our model on multi-organ datasets and then evaluated the performance on six task-specific datasets. Subsequently, we validated the effectiveness of WSI representation learning and conducted comparison experiments with other SOTA methods to showcase the superiority of PAMA. In the end, the ablations and parametric experiments demonstrate the significance of the proposed modules and strategy. Accuracy (ACC), the area under the ROC curve (AUC), and the F1 score were used as metrics to evaluate performance.
@@
Figure 5: Semi-supervised experiments on the Endometrium dataset, utilizing 10%, 35%, 60%, and 85% of labeled data, where fine-tuning results are depicted with solid lines, and linear probing results are denoted with dotted lines.
@@
Table 3: Comparison with SOTA WSI analysis methods on the Endometrium dataset.
@@
4.3 Effectiveness for pan-cancer pre-training
In this experiment, we pre-trained PAMA within the training sets of four datasets, i.e., TCGA-RCC, TCGA-NSCLC, USTC-EGFR, and Endometrium, which are regarded as in-domain datasets. Then, we evaluated the performance of the pre-trained model on the test sets of the in-domain datasets. Additionally, we evaluated its performance on the other two datasets, i.e., TCGA-EGFR and BRCA-HER2, which are regarded as out-of-domain datasets.
For every dataset under DINO features, pre-training on the single dataset can increase ACCs by 0.17% to 22.75%, increase AUCs by 0.31% to 2.15%, and increase F1 score by 0.33% to 19.63% for different tasks. Multi-organ pre-training further promotes the performance of the model. Specifically, on the in-domain datasets, the ACCs/AUCs increase by 29.61%/4.77%, 8.49%/1.85%, 0.81%/1.12%, and 5.84%/2.74% on the Endometrium dataset, TCGA-NSCLC dataset, TCGA-RCC dataset, and USTC-EGFR dataset, respectively. As for the PLIP features, pre-training on the multi-organ datasets can increase ACCs by 1.36% to 17.25%, increase AUCs by 0.93% to 25.28%, and increase F1 score by 1.62% to 17.44% for different tasks. These results demonstrate the proposed method can effectively promote the WSI encoder in optimizing the use of abundant unlabeled WSI data and enhancing representational abilities.
Our model gains even more significant improvement on the Endometrium dataset, where the data is extremely unbalanced. Fig. 4(d)(a) exhibits that the data of LDEA and PECA are less than half of MEDA data, while SDEA, CCEA, and MCEA are even less than ten WSIs. Datasets with long-tailed categories often lead to model bias problems. Fig. 4(d)(d) shows that multi-organ pre-training increased AUCs by 0.04, 0.09, 0.07, 0.27, and 0.12 for categories of LDEA, SDEC, PECA, CCEC, and MCEA, respectively, when compared with the direct training. It demonstrates that PAMA pre-trained on multiple organ datasets can enhance the model generalization ability to significantly relieve the model bias problem.
Molecular characterizations manifest as more latent features that are not visible in histopathology images, and thus molecular status prediction by WSIs is a more challenging task. Pre-training on the single dataset improves the ACCs/F1 score on the TCGA-EGFR dataset and BRCA-HER2 dataset by 1.14%/19.63% and 7.89%/12.09%. It demonstrates that WSI-level self-supervised learning can obtain more discriminative implicit semantic features. Furthermore, directly fine-tuning the multi-organ pre-trained model on the two out-of-domain datasets contributes to an increase in ACCs/F1 scores by 3.43%/24.07% and 13.16%/17.44%. Such a significant improvement indicates that multi-organ pre-training can mine the general semantic information of histopathology images, and thereby can complete various challenging tasks more effectively and efficiently. This demonstrates the ability of our proposed method to be more practical and meaningful in building computer-aided pan-cancer diagnosis systems.
(a) masking ratio $r$
(b) drop out probability $p$
(c) patches per anchor $c$
(d) polar bins $N$
Figure 6: Performance of PAMA as functions of different hyper-parameter settings on the Endometrium validation set.
4.4 Effectiveness of semi-supervised WSI classification
It shows that PAMA is consistently superior to MAE and HIPT across all label ratios. The above results demonstrate the effectiveness of PAMA in pre-training WSI representations and the MIM frameworks are more efficient than the contrastive leaning framework for slide-level learning. PAMA obtains optimal stability in AUCs with label ratios reducing from 85% to 10%. This is of great practical value as it reduces the reliance on a massive number of labeled WSIs for training a robust WSI analysis model. Meanwhile, we can employ unlabeled WSIs with the assistance of PAMA to enhance the capabilities of the WSI analysis models. HIPT is a two-stage pre-training model that is slightly less effective than the one-stage MAE. This illustrates that the discontinuous gradient back-propagation of a multi-stage pre-training model led to an accumulation of biases. In addition, the MAE+ outperforms MAE. It indicates our proposed distance and polar angle embedding can capture more complete spatial information of WSI than the original position encoding of ViT.
@@ 4.5 Comparison with other weakly supervised methods
We compared PAMA with three self-supervised frameworks, BYOL, MAE and HIPT, and four SOTA weakly supervised methods, including DSMIL [25], TransMIL [28], SETMIL [29], and KAT [30] on the Endometrium and TCGA-NSCLC datasets for slide-level classification. The results are shown in Table 4 and 4.
Overall, our proposed PAMA is superior to the second-best method with increased AUCs/ACCs(%) of 0.051/5.49, 0.029/6.05, 0.024/6.18, 0.017/5.87, and 0.013/5.52 on the Endometrium dataset with 10%, 35%, 65%, 85% and 100% labeled data, and increased AUCs/ACCs(%) of 0.027/6.74, 0.017/7.51, 0.015/6.51, 0.012/5.82, and 0.011/4.89 on the TCGA-NSCLC dataset with 10%, 35%, 65%, 85% and 100% labeled training data, respectively.
DSMIL [25] introduced a dual-stream architecture with trainable distance measurement for instances and applied a pyramidal fusion framework for multiscale WSI features, which, however, did not consider the spatial structure of tissue. The absolute structural encoding reduces the performance of DSMIL from other methods. TransMIL and SETMIL leveraged CNN blocks to aggregate local information and then built Transformer structures for long-range global feature aggregation. KAT considered the spatial adjacency of patches and manually defined the fixed hierarchical masks based on local kernels to maintain relative distance information. None of the three methods embedded relative orientation information into slide representations, which causes a significant performance gap compared with PAMA. We re-implemented BYOL [39] with ViT [48] backbone for slide-level feature learning based contrastive learning. Contrastive self-supervised learning frameworks like BYOL rely on extensive and effective augmented views to mine the discriminative representations. However, there are no efficient published WSI-level view augmentation methods currently and we applied random sample patches to construct different views of the WSI. WSI views based on random patch sampling are struggling to efficiently capture semantic information. It results in even worse performance than some weak-supervised methods. Self-supervised learning methods based on MIM, namely, MAE and HIPT, surpass these SOTA weakly supervised methods. It reconfirms the effectiveness of WSI-level representation pre-training.
@@ 4.6 Ablation studies
Table 5: Ablation studies on the Endometrium dataset.
@@
We conducted ablation studies on the Endometrium dataset to verify the significance of our relative spatial embedding and strategy shown in Table. 5. When the polar angel embedding of anchors was removed, we observed the AUC and ACC(%) dropped by 0.016 and 6.39. It is notable that if we applied the polar embedding without the KRO module, the AUC and ACC(%) dropped by 0.029 and 6.87, which means that indexing angles with a fixed polar axis will lead to ambiguous semantic information in WSIs. KRO strategy can dynamically update every anchor’s main polar axis to disambiguate structure information in slides. The relative distance embedding can maintain scale-varying WSIs in a semantic consistency space. The AUC and ACC(%) decreased by 0.022 and 6.75, respectively, when the distance embedding was discarded. When we constructed the slide representation neither with the distance nor polar angle embeddings, the performance had a significant drop of 0.034/7.96 in the AUC/ACC(%). These results prove that the proposed modules can effectively and efficiently acquire spatial information to maintain semantic integrity and consistency in WSIs. Furthermore, the AUC and ACC(%) decreased by 0.004 and 3.83, respectively, when the AD was discarded. This indicates that anchor dropout contributes to better generalization performance.
@@ 4.7 Hyperparameter analysis
To verify the design of the PAMA framework, we performed a series of parametric experiments on the Endometrium dataset, where only the current parameter was tuned in each set of experiments and the remaining parameters were fixed. The results are shown in Fig. 6.
@@
    Polar bins: $N$ is the number of orientation bins, e.g. $N=8$ means each bin holds a $\frac{2\pi}{N}=\frac{\pi}{4}$ angle range. As $N$ increases, the anchor can provide more precise structural information due to the detailed division of orientation intervals. However, this enhancement comes at the cost of increased computational consumption. Based on Fig. 6(d), we set $N=8$ for balancing performance and resource consumption.
@@ 5 Visualization
We further assessed the interpretability of our proposed framework with visualization. We present a well-endometrioid adenocarcinoma slide and the annotation by pathologists as shown in Fig. 7(a-b). Fig. 7(c-f) show the heatmap and polar attention distribution based on anchors in each PACA block without pre-train and during fine-tuning after pre-training with single-organ and multi-organ datasets. In the early stage (e.g. block 1) during fine-tuning after multi-organ pre-training as shown in Fig. 7(III), anchors initially focus on identical pathological tissues as the observation regions. Through supervised WSI label fine-tuning, the anchor’s attention consolidates on high-risk cancerous tissues and attains stability in which the KRO strategy takes a crucial role in adaptively updating the polar axis that is illustrated by the yellow sector in the radar chart. In the process without pre-train as shown in Fig. 7(I), the regions of interest of both anchors in the normal and cancerous tissues are diffuse. After pre-training with the single-organ dataset as shown in Fig. 7(II), the anchor in the positive area can gradually converge to the cancerous tissues. With the contribution of multi-organ pre-training as shown in Fig. 7(III), anchors’ areas of interest are more comprehensive and precise.
@@ 6 Conclusion
In the paper, we focused on self-supervised WSI-level representation learning and proposed the position-aware masked autoencoder (PAMA) for WSI pre-training. The proposed anchor-based position-aware cross-attention (PACA) module leverages the bidirectional communication between the local and global information to capture WSI semantic features. We also introduced a kernel reorientation (KRO) strategy to dynamically update the main orientation of anchors to eliminate ambiguity for WSI representation learning. Additionally, we collected six large-scale datasets from multiple organs and evaluated the effectiveness of PAMA for pan-cancer analysis. The comprehensive experimental results have demonstrated that the proposed method is superior to the state-of-the-art methods and efficiently facilitates the analysis of pan-cancer. The current work has two limitations that can be improved: (1) the collected multi-organ datasets do not yet contain comprehensive cancer types and need to be further expanded for pan-cancer analysis, and (2) the PAMA structure currently relies only on pathology image data, and we need to further introduce multimodal data, such as genomics, to participate in the pre-training process to facilitate cancer diagnosis. In future work, we will focus on these challenges to enhance our work.
Acknowledgments. This work was partly supported by Beijing Natural Science Foundation (Grant No. 7242270), partly supported by the National Natural Science Foundation of China (Grant No. 62171007, 61901018, and 61906058), partly supported by the Fundamental Research Funds for the Central Universities of China (grant No. YWF-23-Q-1075 and JZ2022HGTB0285), partly supported by Emergency Key Program of Guangzhou Laboratory (Grant No. EKPG21-32), partly supported by Joint Fund for Medical Artificial Intelligence (Grant No. MAI2023C014), and partly supported by National Key Research and Development Program of China (Grant No. 2021YFF1201004).

### v4

@@ 4.2 Experimental setting
We first pre-trained our model on multi-organ datasets and then evaluated the performance on six task-specific datasets with two conditions, where the in-domain condition is that fine-tuning datasets are involved in the pre-training, otherwise is the out-of-domain condition. Subsequently, we validated the effectiveness of WSI representation learning and conducted comparison experiments with other SOTA methods to showcase the superiority of PAMA. In the end, the ablations and parametric experiments demonstrate the significance of the proposed modules and strategy. Accuracy (ACC), the area under the ROC curve (AUC), and the F1 score were used as metrics to evaluate performance.
@@
(a) AUCs on various downstream tasks.
(b) ACCs(%) on various downstream tasks.
Figure 5: Performance of different pre-training conditions of PAMA and Prov-GigaPath on various downstream tasks.
4.3 Effectiveness for in-domain pre-training
In this experiment, we pre-trained PAMA within the training sets of five datasets, i.e., TCGA-RCC, TCGA-NSCLC, USTC-EGFR, Endometrium-3k, and TCGA-EGFR, which are regarded as in-domain datasets. Then, we evaluated the performance of the pre-trained model on the test sets of the in-domain datasets to show the effectiveness of PAMA in learning representation from abundant unlabeled histopathology image data.
For every dataset under DINO features, pre-training on the single dataset can increase ACCs by 0.17% to 22.75%, increase AUCs by 0.31% to 2.15%, and increase F1 score by 0.33% to 19.63% for different tasks. Multi-organ pre-training further promotes the performance of the model. Specifically, the ACCs/AUCs increase by 29.61%/4.77%, 8.49%/1.85%, 0.81%/1.12%, 5.84%/2.74%, and 3.43%/4.61% on the Endometrium-3k dataset, TCGA-NSCLC dataset, TCGA-RCC dataset, USTC-EGFR dataset, and TCGA-EGFR dataset, respectively. As for the PLIP features, pre-training on the multi-organ datasets can increase ACCs by 1.36% to 17.25%, increase AUCs by 0.93% to 25.28%, and increase F1 score by 1.62% to 17.44% for different tasks. These results demonstrate the proposed method can effectively promote the WSI encoder in optimizing the use of abundant unlabeled WSI data and enhancing representational abilities.
Our model gains even more significant improvement on the Endometrium-3k dataset, where the data is extremely unbalanced. Fig. 4(a) exhibits that the data of LDEA and PECA are less than half of MEDA data, while SDEA, CCEA, and MCEA are even less than ten WSIs. Datasets with long-tailed categories often lead to model bias problems. Fig. 4(d) shows that multi-organ pre-training increased AUCs by 0.04, 0.09, 0.07, 0.27, and 0.12 for categories of LDEA, SDEC, PECA, CCEC, and MCEA, respectively, when compared with the direct training. It demonstrates that PAMA pre-trained on multiple organ datasets can enhance the model generalization ability to significantly relieve the model bias problem.
Molecular characterizations manifest as more latent features that are not visible in histopathology images, and thus molecular status prediction by WSIs is a more challenging task. Pre-training on the single dataset improves the ACCs/F1 score on the USTC-EGFR dataset and TCGA-EGFR dataset by 3.90%/6.88% and 1.14%/19.63%. It demonstrates that WSI-level self-supervised learning can obtain more discriminative implicit semantic features. Furthermore, directly fine-tuning the multi-organ pre-trained model on the two datasets contributes to an increase in F1 scores by 15.78% and 24.07%. Such a significant improvement indicates that multi-organ pre-training can mine the general semantic information of histopathology images, and thereby can complete various challenging tasks more effectively and efficiently. This demonstrates the ability of our proposed method to be more practical and meaningful in building computer-aided pan-cancer diagnosis systems.
4.4 Generalization for out-of-domain pre-training
We additionally collected a large-scale pan-cancer dataset from TCGA as the out-of-domain data to evaluate the generalization of PAMA pre-training. We pre-trained PAMA and a SOTA method, namely Prov-GigaPath [44], on the pan-cancer dataset without any labels, and then fine-tuned the encoder on six downstream tasks completely independent of the pan-cancer dataset. The results are represented as PAMA (out-of-domain) and Prov-GigaPath (out-of-domain) in Fig. 5. We conducted the experiment using PLIP [20] patch features and the PAMA (in-domain) represents the results of multi-organ strategy in Table 3.
In Fig. 5, the performance of PAMA (out-of-domain) decrease by no more than 0.02 in AUCs and no more than 1.6% in ACCs compared with PAMA (in-domain). Especially on TCGA-NSCLC dataset and TCGA-RCC dataset, the AUCs decreased by less than 0.003, where there is nearly no degradation in performance of pre-training PAMA with the out-of-domain data. The results demonstrate that PAMA exhibits substantial out-of-domain generalization capabilities.
PAMA (out-of-domain) are superior to Prov-GigaPath (out-of-domain) in AUCs by 0.004 to 0.132 and in ACCs by 0.97% to 4.33% for different tasks. It displays a better capacity of our method in characterizing and analyzing unseen data in comparison to Prov-GigaPath.
Furthermore, PAMA (out-of-domain) achieves comparable and superior performance compared with pre-training on the single dataset results in Table 3. The results effectively demonstrate PAMA’s ability to mine information from extensive amounts of unlabeled data, facilitating potential of the framework for more general histopathology image analysis tasks.
@@ 4.4 Generalization for out-of-domain pre-training
Figure 6: Semi-supervised experiments on the Endometrium-3k dataset, utilizing 10%, 35%, 60%, and 85% of labeled data, where fine-tuning results are depicted with solid lines, and linear probing results are denoted with dotted lines.
4.5 Effectiveness of semi-supervised WSI classification
It shows that PAMA is consistently superior to MAE, HIPT and Prov-GigaPath [44] across all label ratios. Prov-GigaPath utilizes DINO V2 [40] to extract patch features and uses LongNet [45] as slide aggregator for pre-training. LongNet was originally designed for extremely long sequences like over 1B+ tokens. However, Prov-GigaPath has not considered any properties of WSI, especially spatial structure information. In sufficient data conditions, it is not even better than HIPT. With the volume of our datasets, the performance of Prov-GigaPath, which is not specifically designed for WSI characteristics, does not differ much from plain MAE, but still has a large margin from PAMA. The above results demonstrate the effectiveness of PAMA in pre-training WSI representations and the MIM frameworks are more efficient than the contrastive leaning framework for slide-level learning. PAMA obtains optimal stability in AUCs with label ratios reducing from 85% to 10%. This is of great practical value as it reduces the reliance on a massive number of labeled WSIs for training a robust WSI analysis model. Meanwhile, we can employ unlabeled WSIs with the assistance of PAMA to enhance the capabilities of the WSI analysis models. HIPT is a two-stage pre-training model that is slightly less effective than the one-stage MAE. This illustrates that the discontinuous gradient back-propagation of a multi-stage pre-training model led to an accumulation of biases. In addition, the MAE+ outperforms MAE. It indicates our proposed distance and polar angle embedding can capture more complete spatial information of WSI than the original position encoding of ViT.
@@ 4.5 Effectiveness of semi-supervised WSI classification
Table 4: Comparison with SOTA WSI analysis methods on the Endometrium-3k dataset.
@@ 4.6 Comparison with other weakly supervised methods
We compared PAMA with four self-supervised frameworks, BYOL, MAE, HIPT, and Prov-GigaPath, and four SOTA weakly supervised methods, including DSMIL [25], TransMIL [28], SETMIL [29], and KAT [30] on the Endometrium-3k and TCGA-NSCLC datasets for slide-level classification. The results are shown in Table 5 and 5.
Overall, our proposed PAMA is superior to the second-best method with increased AUCs/ACCs(%) of 0.051/5.49, 0.029/6.05, 0.024/5.97, 0.017/5.40, and 0.013/5.19 on the Endometrium-3k dataset with 10%, 35%, 65%, 85% and 100% labeled data, and increased AUCs/ACCs(%) of 0.027/6.74, 0.017/6.09, 0.015/5.68, 0.012/5.76, and 0.011/2.83 on the TCGA-NSCLC dataset with 10%, 35%, 65%, 85% and 100% labeled training data, respectively.
DSMIL [25] introduced a dual-stream architecture with trainable distance measurement for instances and applied a pyramidal fusion framework for multiscale WSI features, which, however, did not consider the spatial structure of tissue. The absolute structural encoding reduces the performance of DSMIL from other methods. TransMIL and SETMIL leveraged CNN blocks to aggregate local information and then built Transformer structures for long-range global feature aggregation. KAT considered the spatial adjacency of patches and manually defined the fixed hierarchical masks based on local kernels to maintain relative distance information. None of the three methods embedded relative orientation information into slide representations, which causes a significant performance gap compared with PAMA. We re-implemented BYOL [39] with ViT [53] backbone for slide-level feature learning based contrastive learning. Contrastive self-supervised learning frameworks like BYOL rely on extensive and effective augmented views to mine the discriminative representations. However, there are no efficient published WSI-level view augmentation methods currently and we applied random sample patches to construct different views of the WSI. WSI views based on random patch sampling are struggling to efficiently capture semantic information. It results in even worse performance than some weak-supervised methods. Self-supervised learning methods based on MIM, namely, MAE and Prov-GigaPath, surpass these SOTA weakly supervised methods. It reconfirms the effectiveness of WSI-level representation pre-training.
@@ 4.7 Ablation studies
Table 6: Ablation studies on the Endometrium-3k dataset.
@@
We conducted ablation studies on the Endometrium-3k dataset to verify the significance of our relative spatial embedding and strategy shown in Table. 6. When the polar angel embedding of anchors was removed, we observed the AUC and ACC(%) dropped by 0.016 and 6.39. It is notable that if we applied the polar embedding without the KRO module, the AUC and ACC(%) dropped by 0.029 and 6.87, which means that indexing angles with a fixed polar axis will lead to ambiguous semantic information in WSIs. KRO strategy can dynamically update every anchor’s main polar axis to disambiguate structure information in slides. The relative distance embedding can maintain scale-varying WSIs in a semantic consistency space. The AUC and ACC(%) decreased by 0.022 and 6.75, respectively, when the distance embedding was discarded. When we constructed the slide representation neither with the distance nor polar angle embeddings, the performance had a significant drop of 0.034/7.96 in the AUC/ACC(%). These results prove that the proposed modules can effectively and efficiently acquire spatial information to maintain semantic integrity and consistency in WSIs. Furthermore, the AUC and ACC(%) decreased by 0.004 and 3.83, respectively, when the AD was discarded. This indicates that anchor dropout contributes to better generalization performance.
(a) masking ratio $r$
(b) drop out probability $p$
(c) patches per anchor $c$
(d) polar bins $N$
Figure 7: Performance of PAMA with different hyper-parameter settings on the Endometrium-3k validation set.
@@ 4.8 Hyperparameter analysis
To verify the design of the PAMA framework, we performed a series of parametric experiments on the Endometrium-3k dataset, where only the current parameter was tuned in each set of experiments and the remaining parameters were fixed. The results are shown in Fig. 7.
@@
    Polar bins: $N$ is the number of orientation bins, $e.g.$, $N=8$ means each bin holds a $\frac{2\pi}{N}=\frac{\pi}{4}$ angle range. As $N$ increases, the anchor can provide more precise structural information due to the detailed division of orientation intervals. However, this enhancement comes at the cost of increased computational consumption. Based on Fig. 7(d), we set $N=8$ for balancing performance and resource consumption.
@@ 5 Visualization
We further assessed the interpretability of our proposed framework with visualization. We present a well-endometrioid adenocarcinoma slide and the annotation by pathologists as shown in Fig. 8(a-b). Fig. 8(c-f) show the heatmap and polar attention distribution based on anchors in each PACA block without pre-train and during fine-tuning after pre-training with single-organ and multi-organ datasets. In the early stage ($e.g.$, block 1) during fine-tuning after multi-organ pre-training as shown in Fig. 8(III), anchors initially focus on identical pathological tissues as the observation regions. Through supervised WSI label fine-tuning, the anchor’s attention consolidates on high-risk cancerous tissues and attains stability in which the KRO strategy takes a crucial role in adaptively updating the polar axis that is illustrated by the yellow sector in the radar chart. In the process without pre-train as shown in Fig. 8(I), the regions of interest of both anchors in the normal and cancerous tissues are diffuse. After pre-training with the single-organ dataset as shown in Fig. 8(II), the anchor in the positive area can gradually converge to the cancerous tissues. With the contribution of multi-organ pre-training as shown in Fig. 8(III), anchors’ areas of interest are more comprehensive and precise.
@@
6 Discussion
Most of the current large-scale pathology foundation models focused on patch-level representation learning [20, 21, 22, 54, 41]. A few works focused on slide-level foundation models, such as HIPT [42] and Prov-GigaPath [44], but they disregarded the properties of WSIs, especially the complex spatial semantic information. We introduced the spatial semantic completeness of WSI into the pre-training process, enhancing the slide representations of PAMA to become more semantically complete and generalized.
Data-driven pre-training strategy following foundation models [20, 21, 54, 44, 41] facilitated PAMA for pan-cancer analysis. In this paper, we focused on model design and pan-cancer dataset construction. The proposed position-aware cross-attention model with a dynamical reorientation strategy captures the intrinsic semantic representation of WSIs across various cancer types rather than focusing on any specific tumor or organ. Moreover, the framework was pre-trained to obtain the morphological consistency across multiple cancers based on the broad data including over 13.6k WSIs of 22 cancer types covering 11 organs from multiple medical centers. In future work, it will be necessary to further investigate the spatial properties of pan-cancer and to employ explicit designs to mine its semantic information, such as constructing loss function.
We evaluated the generalization of PAMA across various downstream tasks, including tumor sub-typing, gene mutation prediction, and biomarker status grading. Additionally, out-of-domain datasets were constructed to further demonstrate that the pre-trained model can generalize to datasets not included in the pre-training process. Technically, our pre-training process is task-agnostic, allowing the model to be fine-tuned for specific tasks on any downstream task based on pathology WSIs, which is similar to the released foundation pre-training models [44, 54, 21, 41]. Furthermore, we will explore more general downstream tasks for histopathology image analysis based on the PAMA pre-trained model to facilitate its clinical adoption value.
@@ 7 Conclusion
In the paper, we focused on self-supervised WSI-level representation learning and proposed the position-aware masked autoencoder (PAMA) for WSI pre-training. The proposed anchor-based position-aware cross-attention (PACA) module leverages the bidirectional communication between the local and global information to capture WSI semantic features. We also introduced a kernel reorientation (KRO) strategy to dynamically update the main orientation of anchors to eliminate ambiguity for WSI representation learning. Additionally, we collected seven large-scale datasets from multiple organs and evaluated the effectiveness and generalization of PAMA for pan-cancer analysis. The comprehensive experimental results have demonstrated that the proposed method is superior to the state-of-the-art methods and efficiently facilitates the analysis of pan-cancer. The current work has two limitations that can be improved: (1) the collected multi-organ datasets do not yet contain comprehensive cancer types and need to be further expanded for pan-cancer analysis, and (2) the PAMA structure currently relies only on pathology image data, and we need to further introduce multimodal data, such as genomics, to participate in the pre-training process to facilitate cancer diagnosis. In the future work, we will focus on these challenges to enhance our work.
Acknowledgments. This work was partly supported by Beijing Natural Science Foundation (Grant No. 7242270), partly supported by the National Natural Science Foundation of China (Grant No. 62171007, 61901018, and 61906058), partly supported by the Fundamental Research Fund for the Central Universities of China (grant No. YWF-23-Q-1075), partly supported by the Anhui Provincial Natural Science Foundation (Grant No. 2408085MF162), partly supported by Emergency Key Program of Guangzhou Laboratory (Grant No. EKPG21-32), partly supported by Joint Fund for Medical Artificial Intelligence (Grant No. MAI2023C014), and partly supported by National Key Research and Development Program of China (Grant No. 2021YFF1201004).